OpenStudy (anonymous):

need help 2sqrt5+3sqrt2/5sqrt5+2sqrt2

OpenStudy (studygurl14):

Like this? \(\Large\frac{2\sqrt{5}+3\sqrt{2}}{5\sqrt{5}+2\sqrt{2}}\)

OpenStudy (studygurl14):

Have you heard of "rationalizing the denominator"?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes but tell me the equation to rationalize it

OpenStudy (studygurl14):

Do this: \(\Large\frac{2\sqrt{5}+3\sqrt{2}}{5\sqrt{5}+2\sqrt{2}}\times\frac{5\sqrt{5}-2\sqrt{2}}{5\sqrt{5}-2\sqrt{2}}\)

OpenStudy (studygurl14):

@neoo